Tolu is 5years older than Daniel. Six years ago the product of their ages was 266. How many years from now will the product of their ages be 864
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Tolu is 5years older than Daniel. Six years ago the product of their ages was 266. How many years from now will the product of their ages be 864
-----------
Step 1, find a pair of factors of 266 that differ by 5.
=========
Step 2, find a pair of factors of 864 that differ by 5.

Answer by stanbon(75887)

Tolu is 5years older than Daniel. Six years ago the product of their ages was 266. How many years from now will the product of their ages be 864
------
Equations:
T = D + 5
(T-6)(D-6) = 266
----------------------------
Substitute to solve for T and D::
(D+5-6)(D-6) = 266
(D-1)(D-6) = 266
Solve::
D^2-7D+6 = 266
----
D^2-7D- 260 = 0
(D-20)(D+7) = 0
Positive solution::
D = 20 yrs
So, Tolu is 25 yrs
------------------
Final Equation:
(20+x)(25+x) = 864
x^2 + 45x + 500 = 864
x^2 + 45x - 364 = 0
(x-7)(x+52) = 0
Answer:: 7 years
-----------------------
Cheers,
Stan H.
---------